OpenStudy (anonymous):

Last problem Simplify: 2^3square root x^4 multiplied by ^3 square root 125x^7 I kinda know how to do this now, but the exponents totally lost me

OpenStudy (anonymous):

i got 40x^9 square root of 5x

OpenStudy (ash2326):

is this your question? \[ 2^3 \sqrt{x^4} \sqrt[3]{125x^7} \]

OpenStudy (anonymous):

Yepp

OpenStudy (anonymous):

I just don't know how to make the square root sign.

OpenStudy (ash2326):

ok now \[\sqrt{x^4}= x^2\] \[\sqrt[3]{125x^6*x}=5x^2\sqrt[3]x\] so we get \[8*5*x^4\sqrt[3]{x}\]